Hello,

I need to find the volume of a cylinder wth a 4m radius and a 5m height.

Thanks

Answer by checkley75(3666)

V=B*H BASE=PI*R^2
V=(22/7)*4^2*5
V=(22/7)*16*5
V=1760/7
V=251.42857 M^3